OpenStudy (anonymous):

What is the value of fraction 1 over 3x3 + 5.2y when x = 3 and y = 2? 13.4 16.4 19.4 28.4

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

\[\large \frac{1}{3x^3+5.2y}\] this?

OpenStudy (usukidoll):

is this your equation?

OpenStudy (anonymous):

fraction 1 over 3x3 + 5.2y when x = 3 and y = 2?

OpenStudy (usukidoll):

can you draw your equation?

OpenStudy (anonymous):

it wont work when i draw it

OpenStudy (anonymous):

ya u got it right the first time sry it looked different

OpenStudy (usukidoll):

\(\color{blue}{\text{Originally Posted by}}\) @UsukiDoll \[\large \frac{1}{3x^3+5.2y}\] this? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

so we have x = 3 and y = 2. we just plug it into the equation \[\large \frac{1}{3(3)^3+5.2(2)}\]

OpenStudy (anonymous):

x=3 y=2

OpenStudy (usukidoll):

so what is 5.2 x 2 what is .. 3 x 3 x 3 x 3 ?

OpenStudy (anonymous):

10.4 81

OpenStudy (usukidoll):

\[\large \frac{1}{81+10.4}\] what is 81+10.4 ?

OpenStudy (anonymous):

91.4

OpenStudy (usukidoll):

\[\large \frac{1}{91.4}\] what is 1/91.4

OpenStudy (usukidoll):

it's safe to divide this time. we need the value.

OpenStudy (anonymous):

i dont understand what i have to divide

OpenStudy (anonymous):

r u there?????

OpenStudy (usukidoll):

yeah .. 1 divided by 91.4

OpenStudy (anonymous):

91.4

OpenStudy (usukidoll):

no.

OpenStudy (usukidoll):

it should be a super small decimal

OpenStudy (anonymous):

0.01

OpenStudy (usukidoll):

ya

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but that is not one of the choices

OpenStudy (usukidoll):

what are the choices?

OpenStudy (anonymous):

13.4 16.4 19.4 28.4

OpenStudy (anonymous):

those r all the options

OpenStudy (usukidoll):

then the equation must be off 3(3)+5.2(2) try compute that part only

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

what is the answer

OpenStudy (usukidoll):

what is 3 x 3 what is 5.2 x 2 or provide an accurate screenshot of the equation... otherwise all sorts of errors pop up.

OpenStudy (anonymous):

3x3=9 5.2x2=10.4

OpenStudy (usukidoll):

ok what's 9+10.4

OpenStudy (anonymous):

19.4

OpenStudy (anonymous):

oh i see it

OpenStudy (anonymous):

thx byeeee